OpenStudy (brittany16):

Solve for the variable. -36 + r = -52

OpenStudy (anonymous):

-36+r=-53 r=-53+36 r=-17

jimthompson5910 (jim_thompson5910):

-36 + r = -52 -36 + r + 36 = -52+36 r = -16